# A Survey on Statistical Theory of Deep Learning: Approximation, Training Dynamics, and Generative Models

### v1

@@ 3.2 Limitations of NTK and Beyond the Kernel Regime
3.3 Mean-Field Perspective
A “mean-field” viewpoint is another interesting paradigm to help us understand the optimization landscape of neural network models. Recall that neural network dynamics in the Mean-Field (MF) regime corresponds to $\alpha=1$ in (9). The scaling difference between NTK (i.e., $\alpha=\sqrt{M}$) and MF (i.e., $\alpha=1$) not only leads to entirely distinct neural network behaviors during the training stages of GD or SGD but also influences the scope of viable mathematical analysis. We summarize the differences in comparison Table 2.
The term mean-field comes from an analogy with mean field models in mathematical physics, which analyze the stochastic behavior of many identical particles Ryzhik (2023). Let us denote $\theta_{r}:=(a_{r},w_{r})\in\mathbb{R}^{d+1}$, $\sigma_{\star}(\mathbf{x},\theta_{r}):=a_{r}\sigma(w_{r}^{\top}\mathbf{x})$ in (9). Weight pairs $\{\theta_{r}\}_{r=1}^{M}$ are considered as a collection of gas particles in $\mathcal{D}$-dimensional spaces with $\mathcal{D}:=d+1$. We consider there are infinitely many gas particles allowing $M\rightarrow{\infty}$ which yields the following integral representation of neural dynamics:
$\frac{1}{M}\sum_{r=1}^{M}\sigma_{\star}(\mathbf{x};\mathbf{\theta}_{r})\xrightarrow{M\rightarrow\infty}\int\sigma_{\star}(\mathbf{x};\mathbf{\theta})\rho(d\mathbf{\theta}),$ (17)
Under this setting, the seminal work Mei et al. (2018) studied the evolution of particles $\theta^{(k)}\in\mathbb{R}^{\mathcal{D}}$ updated by $k$-steps of one-pass SGD (training examples are visited once) under $\ell_{2}$-loss. Interestingly, they proved the trajectories of empirical distribution of $\theta^{(k)}$ denoted as $\widehat{\rho}^{(M)}_{k}:=\frac{1}{M}\sum_{r=1}^{M}\delta_{\theta^{(k)}_{r}}$ weakly converges to the deterministic limit $\rho_{t}\in\mathcal{P}(\mathbb{R}^{\mathcal{D}})$ as $k\rightarrow{\infty}$ and $M\rightarrow{\infty}$. The measure $\rho_{t}$ is the solution of the following nonlinear partial differential equation (PDE):
The above PDE describes the evolution of each particle $(\theta_{r})$ in the force field created by the densities of all the other particles. Denote $\mathcal{R}(\rho_{t}):=\mathbb{E}[(y-f(\mathbf{x};\rho_{t}))^{2}]$, then we have
$\sup_{0\leq t\leq T}\big{|}\mathcal{R}(\rho_{t})-\mathcal{R}_{M}(\rho_{\lfloor t/\varepsilon\rfloor})\big{|}\leq e^{C(T+1)}\cdot\sqrt{\frac{1}{M}\vee\varepsilon}\cdot\sqrt{D+\log\frac{M}{\varepsilon}}.$ (19)
The condition for the bound to vanish to 00 is (1) $M\gg D$, (2) $\eta:=\frac{\varepsilon}{2}\ll\frac{1}{D}$, and (3) PDE converges in $T=\mathcal{O}(1)$ iterations. It is interesting to note that the generic ODE approximation requires the step-size $\eta$ to be less than the order of the total number of parameters in the model (i.e., $\eta\ll\frac{1}{MD}$), whereas in this setting the step-size $\eta\ll\frac{1}{D}$ should be enough. Also, recall that the number of sample size $n$ is equivalent to the iteration steps $k:=\lfloor\frac{T}{\varepsilon}\rfloor$ of one-pass SGD with $T=\mathcal{O}(1)$. Then, this means $n=\mathcal{O}(D)\ll\mathcal{O}(MD)$ should be enough for a good approximation. Another notable fact is; in contrast to the kernel regime, the evolution of weights $\theta_{r}$ are non-linear, and in particular, the weights move away from their initialization during training. Indeed under mild assumption, we can show that for small enough step size $\eta$, $\lim_{M\rightarrow\infty}\|\theta^{(k)}-\theta^{(0)}\|_{2}^{2}/M=\Omega(\eta^{2})$ in the mean-field regime, while $\sup_{t\geq 0}\|\theta^{(t)}-\theta^{(0)}\|_{2}^{2}/M=\mathcal{O}(n/(Md))$ in the linear regime. See Bartlett et al. (2021).
NTK Regime MF Regime
Pros
1. Same scaling as in practice
2. Finite time convergence rate
3. Generalization bounds
1. Does not require $\theta^{(k)}$ to be close to $\theta^{(0)}$
2. Potentially learn a larger class of functions
Cons Require $\theta^{(k)}$ to be close to $\theta^{(0)}$
1. Not the same scaling as in practice
2. No finite-time convergence rate
3. No generalization bounds
Table 2: NTK v.s. Mean-Field regimes
Despite the nice characterizations of SGD dynamics through DD, still, the bound in (19) has room for improvement; the number of neurons $M$ is dependent on the ambient data dimension $d$ and the bound is only applicable to the SGD with short convergence iterations $T=\mathcal{O}(1)$. A follow-up work Mei et al. (2019) has attempted to tackle these challenges. Particularly, they proved that there exists a constant $K$ that only depends on intrinsic features of the activation and data distribution, such that with high probability, the following holds:
$\sup_{0\leq t\leq T}\big{|}\mathcal{R}(\rho_{t})-\mathcal{R}_{M}(\rho_{\lfloor t/\varepsilon\rfloor})\big{|}\leq Ke^{K(T\eta)^{3}}\bigg{\{}\sqrt{\frac{\log(M)}{M}}+\sqrt{d+\log(M)}\sqrt{\eta}\bigg{\}}.$ (20)
A remarkable feature of this bound is that as long as $T\eta=\mathcal{O}(1)$ and $K=\mathcal{O}(1)$, the number of neurons only needs to be chosen $M\gg 1$ for the mean-field approximation to be accurate. The condition $T\eta=\mathcal{O}(1)$ mitigates the exponential dependence on $T$ and the bound does not need to scale with the ambient dimension $d$. Later, researchers from the same group generalize the result into multi-layer settings Nguyen and Pham (2020).
3.4 Unifying View of NTK and Mean-Field Regime
In the previous subsections, we have discussed the dynamics of neural networks in two different regimes; NTK and Mean-Field. Despite both being in overparameterized regimes, differences in the scale of dynamics yield two completely different behaviors. Motivated by this observation, there have been attempts to give a unifying view of these two regimes.
@@ 3.4 Unifying View of NTK and Mean-Field Regime
The first attempt we are aware of is Chen et al. (2020b). The motivation of this paper is summarized in the above table, along with the pros and cons of NTK and Mean-Field regimes. As mentioned in Table 2, the caveat of MF analysis lies in its difficulties in obtaining the generalization bound, yet the regime allows the parameters to travel far away from their initializations. The key idea of the paper is to bring in the NTK which requires $\|\theta^{(k)}-\theta^{(0)}\|_{\text{2}}^{2}$ to be small. Instead, they worked on the probability measure space $\mathcal{P}(\mathbb{R}^{d})$ minimizing the following energy functional, where $f(\mathbf{x};\rho):=\alpha\int\sigma_{\star}(\mathbf{x};\mathbf{\theta})\rho(d\mathbf{\theta})$ is the integral representation of neural dynamics in (9):
@@
with a hyper-parameter $\lambda>0$. Here, $\rho^{\star}\in\mathcal{P}(\mathbb{R}^{d})$ is a global minimizer of (21). Note that they penalize the KL-divergence between $\rho$ and $\rho_{0}$ instead the distance between $\theta^{(k)}$ and $\theta^{(0)}$. The training of the objective functional (21) is often implemented through the Wasserstein gradient flow Mei et al. (2018). Note that the noisy gradient descent (NGD) corresponds to a discretized version of the Wasserstein gradient flow in parameter space, and it has been extensively studied in Mei et al. (2018, 2019); Chizat and Bach (2018) that the NGD algorithm approximates the joint measure $\rho^{\star}$ which is the solution of (18) with an additional diffusion term.
@@
This implies that when $\alpha$ in (9) is large, i.e., $\alpha=\mathcal{O}(\sqrt{M})$ (resp. $\alpha=\mathcal{O}(1)$), two-layer neural networks with infinite width trained by noisy gradient descent (NGD) can learn the function class $\mathcal{F}_{\mathcal{X}^{2}}$ (resp. the function class $\mathcal{F}_{\textbf{KL}}$). The two function classes are defined as follows:
@@
Another work Yang and Hu (2020) provides a unifying framework encompassing the parametrizations that induce the neural dynamics in NTK and MF regimes. Under the heavy overparameterized setting (i.e., the width of the network is very large), the main differences between these two regimes come from the different scales on initialized weight parameters and learning rates in GD or SGD. In the paper, the proposed abc-parameterization provides a guideline on how to choose the scales on these parameters. Specifically, recall the definition of a fully connected network with $L+1$ layers in (1). In this context, the width of the network remains consistent at $M$ across all layers. The parametrization is given by a set of numbers $\{a_{\ell},b_{\ell}\}_{\ell}\cup\{c\}$. The pair $(a_{\ell},b_{\ell})$ controls the scale of initialized weight parameters over the layers, and $c$ is the parameter relevant to the learning rate of SGD.
@@
Borrowing the paper’s notation, let us denote $\mathbf{x}_{t}^{\ell}(\mathbf{\xi})$ as the output of the $\ell$-th hidden layer of fully-connected network (1) (after the activation) with the input $\mathbf{\xi}\in\mathbb{R}^{d}$ for the $t$-th SGD updates. The paper defines the difference vector $\Delta\mathbf{x}_{t}^{\ell}(\mathbf{\xi}):=\mathbf{x}_{t}^{\ell}(\mathbf{\xi})-\mathbf{x}_{0}^{\ell}(\mathbf{\xi})$ for all $\ell\in[L+1]$ with $r$ notation:
@@
The corresponding settings of $(a_{\ell},b_{\ell})\cup\{c\}$ for NTK and MF regimes are summarized in the above table. They only focus on the scaling with $M$ and ignore dependence on the intrinsic dimension $d$. The setting for NTK matches with the Theorem $5$ in Du et al. (2019), where the entries of $W_{\ell}$ is drawn from $\mathcal{N}(0,1/M)$ with width-independent $\eta>0$. The reason why we are seeing $a_{\ell}=0$ for $\ell=1$ in the NTK regime is to ensure that the output of every hidden layer in the network is of constant order so that parameterization is in a stable regime. For the MF regime, Mei et al. (2018, 2019) gives us the width-independent parameterizations for weight initializations and learning rate in shallow neural network. Note that the $\frac{1}{M}$-scale in the output weight layer $(\ell=2)$ is canceled with the $M$-scale of learning rate in SGD.

### v3

@@ Appendix D Beyond the Kernel Regime
Appendix E Unifying View of NTK and Mean-Field Regime
@@
There have been attempts to give a unifying view of these two regimes. The first attempt we are aware of is Chen et al. (2020b). The motivation of this paper is summarized in the above table, along with the pros and cons of NTK and Mean-Field regimes. As mentioned in Table 1, the caveat of MF analysis lies in its difficulties in obtaining the generalization bound, yet the regime allows the parameters to travel far away from their initializations. The key idea of the paper is to bring in the NTK which requires $\|\theta^{(k)}-\theta^{(0)}\|_{\text{2}}^{2}$ to be small. Instead, they worked on the probability measure space $\mathcal{P}(\mathbb{R}^{d})$ minimizing the following energy functional, where $f(\mathbf{x};\rho):=\alpha\int\sigma_{\star}(\mathbf{x};\mathbf{\theta})\rho(d\mathbf{\theta})$ is the integral representation of neural dynamics in Equation. (8):
@@
with a hyper-parameter $\lambda>0$. Here, $\rho^{\star}\in\mathcal{P}(\mathbb{R}^{d})$ is a global minimizer of Equation. (32). Note that they penalize the KL-divergence between $\rho$ and $\rho_{0}$ instead the distance between $\theta^{(k)}$ and $\theta^{(0)}$. The training of the objective functional Equation. (32) is often implemented through the Wasserstein gradient flow Mei et al. (2018b). Note that the noisy gradient descent (NGD) corresponds to a discretized version of the Wasserstein gradient flow in parameter space, and it has been extensively studied in Mei et al. (2018b, 2019); Chizat and Bach (2018) that the NGD algorithm approximates the joint measure $\rho^{\star}$ which is the solution of Equation. (28) with an additional diffusion term.
@@
This implies that when $\alpha$ in Equation. 8 (in main text) is large, i.e., $\alpha=\mathcal{O}(\sqrt{M})$ (resp. $\alpha=\mathcal{O}(1)$), two-layer neural networks with infinite width trained by noisy gradient descent (NGD) can learn the function class $\mathcal{F}_{\mathcal{X}^{2}}$ (resp. the function class $\mathcal{F}_{\textbf{KL}}$). The two function classes are defined as follows:
@@
Another work Yang and Hu (2020) provides a unifying framework encompassing the parametrizations that induce the neural dynamics in NTK and MF regimes. Under the heavy overparameterized setting (i.e., the width of the network is very large), the main differences between these two regimes come from the different scales on initialized weight parameters and learning rates in GD or SGD. In the paper, the proposed abc-parameterization provides a guideline on how to choose the scales on these parameters. Specifically, recall the definition of a fully connected network with $L+1$ layers in Equation. 1 (in main text). In this context, the width of the network remains consistent at $M$ across all layers. The parametrization is given by a set of numbers $\{a_{\ell},b_{\ell}\}_{\ell}\cup\{c\}$. The pair $(a_{\ell},b_{\ell})$ controls the scale of initialized weight parameters over the layers, and $c$ is the parameter relevant to the learning rate of SGD.
@@
Borrowing the paper’s notation, let us denote $\mathbf{x}_{t}^{\ell}(\mathbf{\xi})$ as the output of the $\ell$-th hidden layer of fully-connected network Equation. 1 (in main text) (after the activation) with the input $\mathbf{\xi}\in\mathbb{R}^{d}$ for the $t$-th SGD updates. The paper defines the difference vector $\Delta\mathbf{x}_{t}^{\ell}(\mathbf{\xi}):=\mathbf{x}_{t}^{\ell}(\mathbf{\xi})-\mathbf{x}_{0}^{\ell}(\mathbf{\xi})$ for all $\ell\in[L+1]$ with $r$ notation:
@@
The corresponding settings of $(a_{\ell},b_{\ell})\cup\{c\}$ for NTK and MF regimes are summarized in the above table. They only focus on the scaling with $M$ and ignore dependence on the intrinsic dimension $d$. The setting for NTK matches with the Theorem $5$ in Du et al. (2019), where the entries of $W_{\ell}$ is drawn from $\mathcal{N}(0,1/M)$ with width-independent $\eta>0$. The reason why we are seeing $a_{\ell}=0$ for $\ell=1$ in the NTK regime is to ensure that the output of every hidden layer in the network is of constant order so that parameterization is in a stable regime. For the MF regime, Mei et al. (2018b, 2019) gives us the width-independent parameterizations for weight initializations and learning rate in shallow neural network. Note that the $\frac{1}{M}$-scale in the output weight layer $(\ell=2)$ is canceled with the $M$-scale of learning rate in SGD.